# How Primordial Black Holes Change BBN

### v1

@@ 1 INTRODUCTION
There has been a growing interest in the study of Primordial Black Holes (PBHs) within the fields of astrophysics and cosmology. Unlike supermassive black holes, which form from the collapse of massive stars, PBHs originate from primordial density fluctuations and are thus of cosmological origin. The diversity in their possible masses motivates the investigation of alternative formation mechanisms for lower-mass black holes. Formed in the early universe, PBHs may arise through various processes generating primordial inhomogeneities [Carr:2020gox, Green:1997sz, Byrnes:2012yx, Kodama:1982sf, Kuhnel:2015vtw, Deng:2017uwc, Deng:2016vzb].
@@
Although the existence of PBHs remains speculative, their wide mass spectrum and potential abundances make them compelling candidates for dark matter and, as we show here, for altering the cosmic evolution. This makes them valuable probes of both particle physics and the thermal history [Carr:2016drx, Green:2020jor, Carr:2020xqk, Bird:2016dcv, Inomata:2017okj, Clesse:2016vqa, Chisholm:2005vm, Fuller:2017uyd].
@@
Another motivation for studying PBHs is their potential role in the formation of cosmic structures, including galaxies and clusters. Their gravitational influence can attract surrounding matter, facilitating the hierarchical growth of large-scale structures. Theoretical models suggest that PBHs could even serve as seeds of the supermassive black holes observed at galactic centers, merging and accreting mass over cosmic time [Carr:2018rid, Hasinger:2020ptw, Carr:2019kxo, Khlopov:2004sc, Kawasaki:2012kn].
@@
The aforementioned examples of PBH physics firmly reside in the gravity sector. Past work to include PBHs in cosmology has focused on effecting a change in the Hubble expansion rate: another application of gravitational physics. However, PBHs can and will evaporate in the early universe [Carr:2009jm, Keith:2020jww, Kawasaki:2004qu]. In essence, PBH evaporation converts mass energy density into radiation. In this work, we will show how the radiation from PBH decay leads to a change in the entropy history of the universe. A change in entropy can have a profound effect on observable data, namely, the primordial element abundances from Big Bang Nucleosynthesis (BBN). Our work here demonstrates a unique connection between gravitation and nuclear physics.
@@
The outline for the remainder of the paper is as follows. To begin, we give a brief review of the early-universe conditions conducive for PBH formation in Sec. 2. We subsequently calculate the PBH mass fraction $\beta$ using the Green, Liddle, Malik, and Sasaki (GLMS) framework [Green:2004wb] with the inclusion of a running spectral index in Sec. 3. With a model for $\beta$, Sec. 4 details our implementation of Hawking evaporation into a modified version of the burst nuclear-reaction-network code. We use burst to evaluate the impact of PBH decay on the synthesis of the light elements $^{4}{\mathrm{He}}$, $\mathrm{D}$, $^{3}{\mathrm{He}}$, and $^{7}{\mathrm{Li}}$. Sec. 5 gives our computational results, which reveal a distinct transition near PBH mass $M\approx 10^{10}\,\mathrm{g}$. Two distinct regimes of PBH influence on BBN emerge, and we discuss the individual phenomenologies alongside comoving entropy injection from Hawking radiation. We conclude in Sec. 6 with a summary of our findings and a discussion of their broader implications and possible extensions.
@@ 2 THE FORMATION OF PRIMORDIAL BLACK HOLES
The early universe was a hot, dense plasma dominated by relativistic particles, with the radiation energy density determined by the background temperature as $\rho_{\mathrm{rad}}=\frac{\pi^{2}}{30}g_{*}T^{4}$, where $g_{*}$ is the relativistic degrees of freedom. From the Friedmann equations, this energy density can also be expressed as $\rho_{\mathrm{rad}}=\frac{3}{8\pi G4t^{2}}\approx 4.5\times 10^{5}\left(\frac{1\,\mathrm{s}}{t}\right)^{2}\,\mathrm{g}\,\mathrm{cm^{-3}}$. If this radiation density collapsed into a Schwarzschild black hole of radius $r_{\mathrm{s}}$, it would be natural to compare $M_{\mathrm{rad}}=\rho_{\mathrm{rad}}V_{\mathrm{s}}$ with the mass of a Schwarzschild black hole, $M_{\mathrm{s}}=\rho_{\mathrm{s}}V_{\mathrm{s}}$, where $\rho_{\mathrm{s}}=M_{\mathrm{s}}/\frac{4}{3}\pi r_{\mathrm{s}}^{3}\sim 10^{17}\left(M_{\odot}/M_{\mathrm{s}}\right)^{2}\,\mathrm{g}\,\mathrm{cm^{-3}}$. From $M_{\mathrm{rad}}\gtrsim M_{\mathrm{s}}$, one finds that at cosmic time $t$, the mass of a possible Schwarzschild black hole would scale as [Carr:2020gox]
@@
Coincidentally, this Schwarzschild black hole mass is comparable to the Hubble mass in a radiation-dominated era at cosmic time $t$ [Liddle:2000cg, Keith:2020jww]
@@
where $\rho(x)$ is the local energy density on a space-like hypersurface and $\bar{\rho}$ is the cosmic average at a given instance. Early work by Carr and Hawking [Carr:1975qj, Carr:1974nx] showed that an overdense region collapses into a PBH if its density contrast exceeds a threshold $\Delta_{\mathrm{min}}$, but remains below an upper bound $\Delta_{\mathrm{max}}$. The lower threshold, which is determined by the Jeans radius, ensures that the region can overcome its own pressure. The upper bound arises because a region with $\Delta>\Delta_{\mathrm{max}}$ has a local geometry of its space-like hypersurface possessing a positive curvature, which is described by a compact and closed Friedmann universe. Unlike regions with $\Delta_{\mathrm{max}}>\Delta>\Delta_{\mathrm{min}}$, which collapse to form PBHs embedded in the background FRW universe, regions with $\Delta>\Delta_{\mathrm{max}}$ evolve as self-contained universes because their geometry is compact and disconnected from the rest of spacetime.
More recent work has revisited these limits, with [Harada:2013epa, Harada:2004pe] refining the estimates of $\Delta_{\mathrm{max}}$, and [Kopp:2010sh] arguing that such estimates depend strongly on the selected geometry, which is not coordinate-invariant and may not impose robust constraints on physical density fluctuations. Therefore, we do not consider the scenario where $\Delta\gg\Delta_{\mathrm{max}}$ and the resulting acausal regions of spacetime.
A more reliable estimate for the lower threshold has been obtained through hydrodynamical simulations [Evans:1994pj, Niemeyer:1999ak, Hawke:2002rf, Musco:2004ak]. Through the examination of a spherical collapse model in a radiation fluid, these studies suggest that the PBH mass is related to the horizon mass at formation via a critical collapse, $M=K(\Delta-\Delta_{\mathrm{c}})^{\gamma}M_{\mathrm{H}}$, where $K$, $\Delta_{\mathrm{c}}$, and $\gamma$ are constants determined by specific simulation models. A representative set of values, $K=11.9$, $\gamma=0.34$, and $\Delta_{\mathrm{c}}=0.7015$ (Gaussian-curve perturbation), yields $M>0.1M_{\mathrm{H}}$ when $\Delta-\Delta_{\mathrm{c}}$ only slightly exceeds $10^{-6}$ [Niemeyer:1999ak]. This framework provides a more accurate connection between density fluctuations and the expected PBH mass. While larger overdensities give rise to more massive PBHs, the lower threshold of the PBH mass can reasonably be identified with the horizon mass given in (2.2), or at least expected to lie within the same order of magnitude.
@@ 3 THE PRIMORDIAL BLACK HOLE MASS FRACTION
References [Green:2004wb, Young:2014ana] estimate the number density of peaks using the statistical approach introduced in [Bardeen:1985tr], where the number of peaks follows $n_{\mathrm{peaks}}(\nu>\nu_{\mathrm{c}})\propto\exp{(-\nu_{\mathrm{c}}^{2}/2)}$, with $\nu_{\mathrm{c}}^{2}\equiv\Delta_{\mathrm{c}}^{2}/\sigma^{2}_{\Delta}$. Here, $\Delta_{\mathrm{c}}$ represents the threshold density contrast for collapse, and $\sigma_{\Delta}^{2}$ is the variance of the smoothed density field, given by
@@
Here, $W(k,R)$ is the Fourier transformed window function. $\mathcal{P}_{\Delta}(k)$ and $\mathcal{P}_{\mathcal{R}}(k)$ denote the power spectrum of the smoothed density contrast and the primordial curvature perturbation, respectively. From [Planck:2018vyg], the curvature power spectrum is described by $\mathcal{P}_{\mathcal{R}}=A_{\mathcal{R}}\left(k/k_{0}\right)^{n_{s}-1}$, where $A_{\mathcal{R}}\approx 2.1\times 10^{-9}$ is the amplitude of the scalar power spectrum, $n_{s}\approx 0.96$ is the scalar spectral index, and $k_{0}=0.05\,\mathrm{Mpc}^{-1}$ is the pivot scale.
@@
In [Green:2004wb], authors relate $\beta$ to the PBH mass $M\approx M_{\mathrm{H}}$ by connecting the comoving scale $R$ to the horizon mass through $M_{\mathrm{H}}=M_{\mathrm{H,eq}}(k_{\mathrm{eq}}R)^{2}\left(g_{*,\mathrm{eq}}/g_{*}\right)^{1/3}$, where $g_{*,\mathrm{eq}}\approx 3$, $k_{\mathrm{eq}}=0.07\Omega_{\mathrm{m}}h^{2}\,\mathrm{Mpc}^{-1}$, $M_{\mathrm{H,eq}}=1.3\times 10^{49}(\Omega_{\mathrm{m}}h^{2})^{-2}\,\mathrm{g}$, and $\Omega_{\mathrm{m}}h^{2}=0.14$. This mapping expresses the PBH mass in terms of comoving scale and cosmological parameters.
@@
Recent observations, however, allow for a scale-dependent scalar spectral index, commonly referred to as a running spectral index. It is parameterized by $n_{s}(k)=n_{s}+\frac{1}{2}\frac{\mathrm{d}n_{s}}{\mathrm{d}\ln(k)}\ln(\frac{k}{k_{0}})$. Current measurements [Planck:2018vyg] give $n_{s}=0.9641\pm 0.0044$ and $\mathrm{d}n_{s}/\mathrm{d}\ln(k)=-0.0045\pm 0.0067$, evaluated at the pivot scale $k_{0}=0.05\,\mathrm{Mpc}^{-1}$.
@@
The introduction of a running spectral index opens the possibility of producing a significant PBH mass fraction even when $n_{s}$ at CMB scales remains close to unity. To account for this effect in our calculation of $\beta(M)$, we replace the scale-invariant form of the spectrum with its running counterpart. Since no closed analytical expression for the variance (3.1) exists in this case, we rely on numerical evaluation using the burst code [Grohs:2015eua, Grohs:2020xxd].
@@ 4 CHANGE IN BBN FROM PBH DECAY
We have established both the physical conditions under which PBHs can form and the statistical framework that determines their expected mass fraction, $\beta$, at the onset of nucleosynthesis. With these ingredients in hand, we now turn to the central question of how the presence of PBHs modifies the dynamics of BBN. By emitting energetic particles through Hawking radiation, PBHs can inject entropy and alter reaction rates in the early plasma, thereby modifying the abundances of light elements relative to standard BBN predictions. Before addressing these modifications, it is useful to briefly recall the main features of standard BBN in the absence of PBHs.
@@
Because of the Coulomb barrier, the synthesis of nuclei beyond $\mathrm{A}=8$ is strongly suppressed, leaving only trace amounts of heavier elements (e.g., $^{9}\mathrm{Be}$, $^{10}\mathrm{B}$). As a result, the observable relics of BBN are $^{1}\mathrm{H}$, $\mathrm{D}$, $^{3}{\mathrm{He}}$, $^{4}{\mathrm{He}}$, also with $^{6}\mathrm{Li}$ and $^{7}{\mathrm{Li}}$ produced at much lower levels. The primordial abundance of $^{6}\mathrm{Li}$ remains uncertain because it is readily destroyed in stellar environments, whereas $^{7}{\mathrm{Li}}$ is consistently overproduced in standard BBN compared with observational data — a long-standing discrepancy known as the “lithium problem” [Jedamzik:2007cp, Asplund:2005yt]. Comparisons between the observed and predicted primordial abundances of these light elements provide valuable insights into the physics of the early universe. In this work, we will focus on $^{4}{\mathrm{He}}$, $\mathrm{D}$, $^{3}{\mathrm{He}}$, and $^{7}{\mathrm{Li}}$.
In standard BBN, nearly all free neutrons are incorporated into $^{4}{\mathrm{He}}$. Consequently, the primordial $^{4}{\mathrm{He}}$ abundance is determined primarily by the neutron-to-proton ratio prior to nucleosynthesis. This ratio can be approximated by the Boltzmann factor $n/p\approx e^{-\Delta m/T_{\mathrm{f}}}\approx 1/5$, where $\Delta m=m_{n}-m_{p}$ and $T_{\mathrm{f}}\approx 0.8\,\mathrm{MeV}$ is the freeze-out temperature [Kolb:1990vq]. Below this temperature, weak interactions involving free neutron decay and lepton capture processes continue to reduce the ratio to $n/p\approx 1/7$, yielding a helium mass fraction of $Y_{\mathrm{P}}=n(^{4}{\mathrm{He}})m(^{4}{\mathrm{He}})/n_{\mathrm{b}}m_{\mathrm{b}}=2(n/p)/[1+(n/p)]\approx 0.25$. Observational determinations of $^{4}{\mathrm{He}}$ from emission lines in H-II regions suggest $Y_{\mathrm{P}}=0.2345\pm 0.0026$ [Peimbert:2000yj], while more recent analyses using Markov Chain Monte Carlo techniques yield $Y_{\mathrm{P}}=0.2449\pm 0.0040$ [Aver:2015iza].
@@
Among the light elements, deuterium ($\mathrm{D}$) is particularly valuable as a probe of the baryon-to-photon ratio, since the deuterium-to-hydrogen ratio ($\mathrm{D}/\mathrm{H}$) decreases monotonically with $\eta$ and is highly sensitive to it [Tytler:2000qf]. A higher baryon density allows nuclear reactions to proceed more efficiently, synthesizing nearly all of the deuterium into helium. Moreover, D/H can be directly measured from absorption spectra in metal-poor, damped Lyman-$\alpha$ systems. Recent observations report $\mathrm{D}/\mathrm{H}=(2.527\pm 0.030)\times 10^{-5}$ [Cooke:2017cwo]. $^{3}{\mathrm{He}}$ provides another sensitive indicator of $\eta$, though its interpretation is complicated by continuous production and destruction in stellar environments [Kawasaki:2017bqm]. As a result, present-day measurements of $^{3}{\mathrm{He}}$ are less robust indicators of $\eta$ than deuterium. Measurements in Galactic H-II regions yield $^{3}{\mathrm{He}}/\mathrm{H}=(1.1\pm 0.2)\times 10^{-5}$ [Bania:2002yj]. $^{7}{\mathrm{Li}}$ remains the most puzzling case: standard BBN, calibrated by the baryon density inferred from CMB data, predicts abundances significantly higher than those observed in metal-poor halo stars, where $^{7}{\mathrm{Li}}/\mathrm{H}=(1.23^{+0.68}_{-0.32})\times 10^{-10}$ [Ryan:1999rxn, Fields:2011zzb].
Having outlined the key features of standard BBN, we now turn to how this picture is modified in the presence of PBHs. PBHs can influence BBN in several ways, but one of the crucial effects arises from Hawking radiation, through which they inject high-energy particles into the primordial plasma. A black hole of mass $M$ radiates particles at a characteristic temperature [Hawking:1975vcx]
@@
Hawking radiation causes the black hole to lose mass at a rate [MacGibbon:1991tj]
@@
The function $f(M)$ accounts for the evaporation rates of different particle species, including leptons $l$, quarks $q$, neutrinos $a$, gluons $g$, $W$ and $Z$ bosons, as well as the Higgs boson. Its explicit form, incorporating spin-dependent weights and degrees of freedom, is given by [Lunardini:2019zob]
@@
where $M_{j}=1/(8\pi Gm_{j})$ corresponds to the PBH mass whose Hawking temperature equals the rest mass $m_{j}$ of a given particle species. Massless photons contribute to the first term, $2f_{1}$, while massless gluons are assigned an effective mass $m_{g}=0.6\,\mathrm{GeV}$, corresponding to the infrared cutoff in QCD renormalization [Halzen:1984mc]. The coefficients $s_{s},f_{s}^{q}$, and $\eta_{\nu}^{N}$ encode the spin and charge dependent parameters, and distinguish between Dirac and Majorana neutrinos. Their numerical values are summarized in

### v2

# arXiv:2511.18646v2[astro-ph.CO] 01 Mar 2026

@@ 1 INTRODUCTION
There has been a growing interest in the study of Primordial Black Holes (PBHs) within the fields of astrophysics and cosmology. Unlike supermassive black holes, which form from the collapse of massive stars, PBHs originate from primordial density fluctuations and are thus of cosmological origin. The diversity in their possible masses motivates the investigation of alternative formation mechanisms for lower-mass black holes. Formed in the early universe, PBHs may arise through various processes generating primordial inhomogeneities [18, 36, 11, 59, 62, 29, 28, 41, 76, 42, 53, 51, 70, 12, 25, 26].
@@
Although the existence of PBHs remains speculative, their wide mass spectrum and potential abundances make them compelling candidates for dark matter and, as we show here, for altering the cosmic evolution. This makes them valuable probes of both particle physics and the thermal history [19, 34, 20, 8, 48, 23, 22, 33, 75, 9].
@@
Another motivation for studying PBHs is their potential role in the formation of cosmic structures, including galaxies and clusters. Their gravitational influence can attract surrounding matter, facilitating the hierarchical growth of large-scale structures. Theoretical models suggest that PBHs could even serve as seeds of the supermassive black holes observed at galactic centers, merging and accreting mass over cosmic time [21, 44, 14, 58, 56].
@@
The aforementioned examples of PBH physics firmly reside in the gravity sector. Past work to include PBHs in cosmology has focused on inducing a change in the Hubble expansion rate: another application of gravitational physics. However, PBHs can and will evaporate in the early universe [13, 57, 55]. In essence, PBH evaporation converts mass energy density into radiation. In this work, we will show how the radiation from PBH decay leads to a change in the entropy history of the universe. A change in entropy can have a profound effect on observable data, namely, the primordial element abundances from Big Bang Nucleosynthesis (BBN). Our work here demonstrates a unique connection between gravitation and nuclear physics.
@@
The outline for the remainder of the paper is as follows. To begin, we give a brief review of the early-universe conditions conducive for PBH formation in Sec. 2. We subsequently calculate the PBH mass fraction $\beta$ using the Green, Liddle, Malik, and Sasaki (GLMS) framework [35] with the inclusion of a running spectral index in Sec. 3. With a model for $\beta$, Sec. 4 details our implementation of Hawking evaporation into a modified version of the burst nuclear-reaction-network code. We use burst to evaluate the impact of PBH decay on the synthesis of the light elements $^{4}{\mathrm{He}}$, $\mathrm{D}$, $^{3}{\mathrm{He}}$, and $^{7}{\mathrm{Li}}$. Sec. 5 gives our computational results, which reveal a distinct transition near PBH mass $M\approx 10^{10}\,\mathrm{g}$. Two distinct regimes of PBH influence on BBN emerge, and we discuss the individual phenomenologies alongside comoving entropy injection from Hawking radiation. We conclude in Sec. 6 with a summary of our findings and a discussion of their broader implications and possible extensions.
@@ 2 THE FORMATION OF PRIMORDIAL BLACK HOLES
There is a wide variety of models for the formation of PBH’s. Below, we focus on a representative model, as an illustration from which we can estimate the mass spectrum, needed for our purposes. The early universe was a hot, dense plasma dominated by relativistic particles, with the radiation energy density determined by the background temperature as $\rho_{\mathrm{rad}}=\frac{\pi^{2}}{30}g_{*}T^{4}$, where $g_{*}$ is the relativistic degrees of freedom. From the Friedmann equations, the energy density can also be expressed as $\rho_{\mathrm{rad}}=\frac{3}{8\pi G4t^{2}}\approx 4.5\times 10^{5}\left(\frac{1\,\mathrm{s}}{t}\right)^{2}\,\mathrm{g}\,\mathrm{cm^{-3}}$. If this radiation density collapsed into a Schwarzschild black hole of radius $r_{\mathrm{s}}$, it would be natural to compare $M_{\mathrm{rad}}=\rho_{\mathrm{rad}}V_{\mathrm{s}}$ with the mass of a Schwarzschild black hole, $M_{\mathrm{s}}=\rho_{\mathrm{s}}V_{\mathrm{s}}$, where $\rho_{\mathrm{s}}=M_{\mathrm{s}}/\frac{4}{3}\pi r_{\mathrm{s}}^{3}\sim 10^{17}\left(M_{\odot}/M_{\mathrm{s}}\right)^{2}\,\mathrm{g}\,\mathrm{cm^{-3}}$. From $M_{\mathrm{rad}}\gtrsim M_{\mathrm{s}}$, one finds that at cosmic time $t$, the mass of a possible Schwarzschild black hole would scale as [18]
@@
Coincidentally, this Schwarzschild black hole mass is comparable to the Hubble mass in a radiation-dominated era at cosmic time $t$ [63, 57]
@@
where $\rho(x)$ is the local energy density on a space-like hypersurface and $\bar{\rho}$ is the cosmic average at a given instance. Early work by Carr and Hawking [16, 15] showed that an overdense region collapses into a PBH if its density contrast exceeds a threshold $\Delta_{\mathrm{min}}$, but remains below an upper bound $\Delta_{\mathrm{max}}$. The lower threshold, which is determined by the Jeans radius, ensures that the region can overcome its own pressure. The upper bound arises because a region with $\Delta>\Delta_{\mathrm{max}}$ has a local geometry of its space-like hypersurface possessing a positive curvature, which is described by a compact and closed Friedmann universe. Unlike regions with $\Delta_{\mathrm{max}}>\Delta>\Delta_{\mathrm{min}}$, which collapse to form PBHs embedded in the background FRW universe, regions with $\Delta>\Delta_{\mathrm{max}}$ evolve as self-contained universes because their geometry is compact and disconnected from the rest of spacetime.
More recent work has revisited these limits, with [43, 40] refining the estimates of $\Delta_{\mathrm{max}}$, and [61] arguing that such estimates depend strongly on the selected geometry, which is not coordinate-invariant and may not impose robust constraints on physical density fluctuations. Therefore, we do not consider the scenario where $\Delta\gg\Delta_{\mathrm{max}}$ and the resulting acausal regions of spacetime.
A more reliable estimate for the lower threshold has been obtained through hydrodynamical simulations [31, 68, 45, 67]. Through the examination of a spherical collapse model in a radiation fluid, these studies suggest that the PBH mass is related to the horizon mass at formation via a critical collapse, $M=K(\Delta-\Delta_{\mathrm{c}})^{\gamma}M_{\mathrm{H}}$, where $K$, $\Delta_{\mathrm{c}}$, and $\gamma$ are constants determined by specific simulation models. A representative set of values, $K=11.9$, $\gamma=0.34$, and $\Delta_{\mathrm{c}}=0.7015$ (Gaussian-curve perturbation), yields $M>0.1M_{\mathrm{H}}$ when $\Delta-\Delta_{\mathrm{c}}$ only slightly exceeds $10^{-6}$ [68]. This framework provides a more accurate connection between density fluctuations and the expected PBH mass. While larger overdensities give rise to more massive PBHs, the lower threshold of the PBH mass can reasonably be identified with the horizon mass given in (2.2), or at least expected to lie within the same order of magnitude.
@@ 3 THE PRIMORDIAL BLACK HOLE MASS FRACTION
References [35, 77] estimate the number density of peaks using the statistical approach introduced in [6], where the number of peaks follows $n_{\mathrm{peaks}}(\nu>\nu_{\mathrm{c}})\propto\exp{(-\nu_{\mathrm{c}}^{2}/2)}$, with $\nu_{\mathrm{c}}^{2}\equiv\Delta_{\mathrm{c}}^{2}/\sigma^{2}_{\Delta}$. Here, $\Delta_{\mathrm{c}}$ represents the threshold density contrast for collapse, and $\sigma_{\Delta}^{2}$ is the variance of the smoothed density field, given by
@@
Here, $W(k,R)$ is the Fourier transformed window function. $\mathcal{P}_{\Delta}(k)$ and $\mathcal{P}_{\mathcal{R}}(k)$ denote the power spectrum of the smoothed density contrast and the primordial curvature perturbation, respectively. From [2], the curvature power spectrum is described by $\mathcal{P}_{\mathcal{R}}=A_{\mathcal{R}}\left(k/k_{0}\right)^{n_{s}-1}$, where $A_{\mathcal{R}}\approx 2.1\times 10^{-9}$ is the amplitude of the scalar power spectrum, $n_{s}\approx 0.96$ is the scalar spectral index, and $k_{0}=0.05\,\mathrm{Mpc}^{-1}$ is the pivot scale.
@@
In [35], authors relate $\beta$ to the PBH mass $M\approx M_{\mathrm{H}}$ by connecting the comoving scale $R$ to the horizon mass through $M_{\mathrm{H}}=M_{\mathrm{H,eq}}(k_{\mathrm{eq}}R)^{2}\left(g_{*,\mathrm{eq}}/g_{*}\right)^{1/3}$, where $g_{*,\mathrm{eq}}\approx 3$, $k_{\mathrm{eq}}=0.07\Omega_{\mathrm{m}}h^{2}\,\mathrm{Mpc}^{-1}$, $M_{\mathrm{H,eq}}=1.3\times 10^{49}(\Omega_{\mathrm{m}}h^{2})^{-2}\,\mathrm{g}$, and $\Omega_{\mathrm{m}}h^{2}=0.14$. This mapping expresses the PBH mass in terms of comoving scale and cosmological parameters.
@@
Recent observations, however, allow for a scale-dependent scalar spectral index, commonly referred to as a running spectral index. It is parameterized by $n_{s}(k)=n_{s}+\frac{1}{2}\frac{\mathrm{d}n_{s}}{\mathrm{d}\ln(k)}\ln(\frac{k}{k_{0}})$. Current measurements [2] give $n_{s}=0.9641\pm 0.0044$ and $\mathrm{d}n_{s}/\mathrm{d}\ln(k)=-0.0045\pm 0.0067$, evaluated at the pivot scale $k_{0}=0.05\,\mathrm{Mpc}^{-1}$.
@@
The introduction of a running spectral index opens the possibility of producing a significant PBH mass fraction even when $n_{s}$ at CMB scales remains close to unity. To account for this effect in our calculation of $\beta(M)$, we replace the scale-invariant form of the spectrum with its running counterpart. Since no closed analytical expression for the variance (3.1) exists in this case, we rely on numerical evaluation using the burst code [37, 38].
@@ 4 CHANGE IN BBN FROM PBH DECAY
We have established both the physical conditions under which PBHs can form and the statistical framework that determines their expected mass fraction, $\beta$, at the onset of nucleosynthesis. With these ingredients in hand, we now turn to the central question of how the presence of PBHs modifies the dynamics of BBN. By emitting energetic particles through Hawking radiation, PBHs can inject entropy and alter reaction rates in the early plasma [74, 73, 9], thereby modifying the abundances of light elements relative to standard BBN predictions. Before addressing these modifications, it is useful to briefly recall the main features of standard BBN in the absence of PBHs.
@@
Because of the Coulomb barrier, the synthesis of nuclei beyond $\mathrm{A}=8$ is strongly suppressed, leaving only trace amounts of heavier elements (e.g., $^{9}\mathrm{Be}$, $^{10}\mathrm{B}$). As a result, the observable relics of BBN are $^{1}\mathrm{H}$, $\mathrm{D}$, $^{3}{\mathrm{He}}$, $^{4}{\mathrm{He}}$, also with $^{6}\mathrm{Li}$ and $^{7}{\mathrm{Li}}$ produced at much lower levels. The primordial abundance of $^{6}\mathrm{Li}$ remains uncertain because it is readily destroyed in stellar environments, whereas $^{7}{\mathrm{Li}}$ is consistently overproduced in standard BBN compared with observational data — a long-standing discrepancy known as the “lithium problem” [52, 3, 50]. Comparisons between the observed and predicted primordial abundances of these light elements provide valuable insights into the physics of the early universe. In this work, we will focus on $^{4}{\mathrm{He}}$, $\mathrm{D}$, $^{3}{\mathrm{He}}$, and $^{7}{\mathrm{Li}}$.
In standard BBN, nearly all free neutrons are incorporated into $^{4}{\mathrm{He}}$. Consequently, the primordial $^{4}{\mathrm{He}}$ abundance is determined primarily by the neutron-to-proton ratio prior to nucleosynthesis. This ratio can be approximated by the Boltzmann factor $n/p\approx e^{-\Delta m/T_{\mathrm{f}}}\approx 1/5$, where $\Delta m=m_{n}-m_{p}$ and $T_{\mathrm{f}}\approx 0.8\,\mathrm{MeV}$ is the freeze-out temperature [60]. Below this temperature, weak interactions involving free neutron decay and lepton capture processes continue to reduce the ratio to $n/p\approx 1/7$, yielding a helium mass fraction of $Y_{\mathrm{P}}=n(^{4}{\mathrm{He}})m(^{4}{\mathrm{He}})/n_{\mathrm{b}}m_{\mathrm{b}}=2(n/p)/[1+(n/p)]\approx 0.25$. Observational determinations of $^{4}{\mathrm{He}}$ from emission lines in H-II regions suggest $Y_{\mathrm{P}}=0.2345\pm 0.0026$ [69], while more recent analyses using Markov Chain Monte Carlo techniques yield $Y_{\mathrm{P}}=0.2449\pm 0.0040$ [4].
@@
Among the light elements, deuterium ($\mathrm{D}$) is particularly valuable as a probe of the baryon-to-photon ratio, since the deuterium-to-hydrogen ratio ($\mathrm{D}/\mathrm{H}$) decreases monotonically with $\eta$ and is highly sensitive to it [72, 49]. A higher baryon density allows nuclear reactions to proceed more efficiently, synthesizing nearly all of the deuterium into helium. Moreover, D/H can be directly measured from absorption spectra in metal-poor, damped Lyman-$\alpha$ systems. Recent observations report $\mathrm{D}/\mathrm{H}=(2.527\pm 0.030)\times 10^{-5}$ [24]. $^{3}{\mathrm{He}}$ provides another sensitive indicator of $\eta$, though its interpretation is complicated by continuous production and destruction in stellar environments [54]. As a result, present-day measurements of $^{3}{\mathrm{He}}$ are less robust indicators of $\eta$ than deuterium. Measurements in Galactic H-II regions yield $^{3}{\mathrm{He}}/\mathrm{H}=(1.1\pm 0.2)\times 10^{-5}$ [5]. $^{7}{\mathrm{Li}}$ remains the most puzzling case: standard BBN, calibrated by the baryon density inferred from CMB data, predicts abundances significantly higher than those observed in metal-poor halo stars, where $^{7}{\mathrm{Li}}/\mathrm{H}=(1.23^{+0.68}_{-0.32})\times 10^{-10}$ [71, 32].
Having outlined the key features of standard BBN, we now turn to how this picture is modified in the presence of PBHs. PBHs can influence BBN in several ways, but one of the crucial effects arises from Hawking radiation, through which they inject high-energy particles into the primordial plasma. A black hole of mass $M$ radiates particles at a characteristic temperature [46]
@@
Hawking radiation causes the black hole to lose mass at a rate [65]
@@
The function $f(M)$ accounts for the evaporation rates of different particle species, including leptons $l$, quarks $q$, neutrinos $a$, gluons $g$, $W$ and $Z$ bosons, as well as the Higgs boson. Its explicit form, incorporating spin-dependent weights and degrees of freedom, is given by [64]
@@
where $M_{j}=1/(8\pi Gm_{j})$ corresponds to the PBH mass whose Hawking temperature equals the rest mass $m_{j}$ of a given particle species. Massless photons contribute to the first term, $2f_{1}$, while massless gluons are assigned an effective mass $m_{g}=0.6\,\mathrm{GeV}$, corresponding to the infrared cutoff in QCD renormalization [39]. The coefficients $s_{s},f_{s}^{q}$, and $\eta_{\nu}^{N}$ encode the spin and charge dependent parameters, and distinguish between Dirac and Majorana neutrinos. Their numerical values are summarized in